OpenStudy (anonymous):

When two bodies , a heavy one and a light one are acted upon by the same force for the same time then :- a)the light body builds up a lower velocity than the heavy one and the momentum gained is different in both the cases. b)the light body builds up a higher velocity than the heavy one and the momentum gained is different in both the cases. c)the light body builds up a higher velocity than the heavy one and the momentum gained is same in both the cases. Please justify your answer also !!

OpenStudy (abhisar):

Force = Rate of change of momentum or \(\sf Force = \huge \frac{\Delta mv}{t}\)

OpenStudy (abhisar):

Here Force and time are constant

OpenStudy (nincompoop):

F = ma

OpenStudy (anonymous):

the force of gravity will be differnt on both the objects ? the heavier one will accelerate faster ?

OpenStudy (abhisar):

so \(\sf m\propto \frac{1}{v}\)

OpenStudy (abhisar):

Heavier one will have less velocity

OpenStudy (nincompoop):



OpenStudy (anonymous):

heavier one will hv less velocity so the ans is either b or c ,now what about momentum ?

OpenStudy (abhisar):

and since the force is same for both, momentum will be equal

OpenStudy (nincompoop):

you can reconcile this even with simple F = ma where acceleration is the change in velocity over time. heavier objects tend to require more force to keep up with the speed to keep up with lighter objects

OpenStudy (nincompoop):

momentum is a separate issue

OpenStudy (anonymous):

i cannot understand why momentum is same ?

OpenStudy (abhisar):

Because Force = Rate of change in momentum. Same force for same time will build up same momentum

OpenStudy (nincompoop):

momentum of a lighter objects that speeds up really really fast may be the same for heavier objects with slower speed it is based on the formula you were given earlier

OpenStudy (abhisar):

and since momentum = mass*velocity, the one with lower mass will have higher velocity and vice-versa

OpenStudy (anonymous):

so u mean to say tht the force will be same on both the heavy and ligt objects ?

OpenStudy (abhisar):

yes

OpenStudy (anonymous):

\[Ma = m a _{1}\] and M > m then a1 > a

OpenStudy (nincompoop):

momentum (P) P = mass * velocity

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

but see it https://www.youtube.com/watch?v=_mCC-68LyZM it says tht force is different !

OpenStudy (nincompoop):

you need to understand kinetic and potential energies to get a better clue how momentum works

OpenStudy (anonymous):

you can think of it without momentum

OpenStudy (anonymous):

see 1:32 section of this video

OpenStudy (nincompoop):

http://finedrafts.com/files/CUNY/Physics/ you can select either physics with calculus or just with algebra, the concept remains the same

OpenStudy (nincompoop):

stop watching videos and start reading textbooks first

OpenStudy (anonymous):

ohk :)

OpenStudy (abhisar):

I think it would be better if one person tries to make him understand at once.

OpenStudy (anonymous):

yes please i am really confused about momentum stuff here

OpenStudy (anonymous):

this video is saying tht they accelerate at differnt rate !

OpenStudy (anonymous):

choices tell you about building up means acceleration it is the same force he told you that and that video spoke about gravitation force !!

OpenStudy (nincompoop):

start with textbooks then we can clarify them all later start with speed/velocity, acceleration ... it is a progression of concepts

OpenStudy (nincompoop):

acceleration at different rates has no direct bearing on momentum because momentum depends on both the mass and velocity two factors determine it!

OpenStudy (anonymous):

ohk

OpenStudy (nincompoop):

it's why you need to understand other concepts like KINETIC and POTENTIAL ENERGIES first! you will see how they beautifully fit into momentum

OpenStudy (abhisar):

@moli1993 Force is equal to rate of change in momentum i.e Force = \(\sf \Delta mv/t\) is this ok ?

OpenStudy (abhisar):

\(\sf \Delta mv\) means Final momentum-initial momentum

OpenStudy (abhisar):

So, Final momentum - initial momentum = Force*time ok ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[=m \Delta v \]

OpenStudy (abhisar):

Now suppose a body is initially at rest

OpenStudy (abhisar):

This will lead to 0 initial momentum....correct ?

OpenStudy (anonymous):

yes

OpenStudy (abhisar):

Now suppose F force for time t is applied on it. then its momentum will become F*T and this will be equal to product of its mass and velocity...correct ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but f*t = impulse ?

OpenStudy (abhisar):

If we change the mass of this object [somehow] then its velocity will increase accordingly to maintain the product equal to F*t

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @moli1993 but f*t = impulse ? \(\color{blue}{\text{End of Quote}}\) Yes that's correct. you can also say that change in momentum= Impulse

OpenStudy (anonymous):

why should velocity increase to maintain f*t ??

OpenStudy (abhisar):

Because the product F*t has to be constant. that's the condition we are dealing with

OpenStudy (anonymous):

ok ok then ?

OpenStudy (nincompoop):

\[J = \sum F (t_2 - t_1) = \sum F \Delta t\] \[\sum F = \frac{p_2-p_1}{t_2-t_1} \rightarrow \sum F(t_2-t_1) = p_2-p_1 \rightarrow J = p_2 - p_1 \] Impulse-momentum theorem

OpenStudy (abhisar):

So basically, whet F*t is constant, m is inversely proportional to velocity

OpenStudy (abhisar):

in other words, larger mass will have lower velocity and smaller mass will have larger velocity

OpenStudy (abhisar):

and this is because the product F*t is constant which means momentum is constant. Does that make any sense to you ?

OpenStudy (nincompoop):

Impulse when force is not constant \[\int\limits^{t_2}_{t_1} \sum F dt = \int\limits^{t_2}_{t_1}\frac{dp}{dt}dt = \int\limits^{p_2}_{p_1} dp = p_2 - p_1\] general definition of impluse \[J = \int\limits^{t_2}_{t_1} \sum F dt\]

OpenStudy (nincompoop):

do you get it yet? one way to reconcile this is to solve an actual problem with varying mass and same force applied. calculate the momentum, but first you need to calculate the velocity

OpenStudy (anonymous):

Thank u @nincompoop and @Abhisar :)